OpenStudy (anonymous):

4ab(squared)-3ab(squared) directions: simplify the expression using only positive exponents in your answer

OpenStudy (anonymous):

I guess you mean \[4a(b^2) - 3a(b^2)\]right? If so, this simplifies to \[a(b^2)\]

OpenStudy (anonymous):

thats right, but i just dont know how to do squared on the computer lol but is there anyway you could possibly explain to me how you got that, so i can understand better>

OpenStudy (anonymous):

Yep, you can think of \[a(b^2)\]as being an item on its own, so you're taking 3 of those items from 4 of them. Algebraically this can be written as: \[4*a(b^2)-3*a(b^2) = (4-3)*(a(b^2))\]Which equals: \[(1)*a(b^2) = a(b^2)\]Hopefully that clears things up a bit.

OpenStudy (anonymous):

it definately did. thanks!

OpenStudy (anonymous):

No problem!